# Label Set Optimization via Activation Distribution Kurtosis for Zero-shot Classification with Generative Models

### v1

@@ 4.3 Suggestions to Practitioners
    The optimal lexical choices for label options is both dataset-dependent and model-dependent.
  4. 4.
@@ 5 Neuron Analysis for Label Selection
Although our findings indicate the importance of label name selection for classification in zero-shot ICL, current studies lack consideration of this factor. Therefore, we conduct analysis to gain insights into the underlying mechanism of lexical choice for single-word label names. Inspired by studies on pre-trained language models or LLMs that indicate the correlation between neuron activation pattern in FFN and model performance (Kuzmin et al., 2023; Tang et al., 2024; Stolfo et al., 2024; Wu et al., 2024), we hypothesise the model performance influenced by label names may correlate with the number of outliers in the neurons within FFN in the decoder of the LLMs.
@@
where $\tilde{h}^{i}$ is the output hidden states from multi-head self-attention module. The activation function (act_fn) for Flan-T5 and Llama 3/3.1 is Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel, 2016) and Sigmoid Linear Unit (SiLU) (Hendrycks and Gimpel, 2016; Elfwing et al., 2018), respectively.
[…]
@@
where $\mu$ is the mean of $N_{I}$. For each dataset, we average the kurtosis score over the validation set for each candidate label set. We then calculate the Spearman correlation between model performance and the averaged kurtosis score.
@@ Results
Figure 3: Macro-F1 vs. neuron activation kurtosis for the emergent dataset with different label options in the prompt for Llama 3.
Table 4 shows that, for most datasets, there is statistically significant negative correlation between model performances and kurtosis scores across models and activation functions, indicating that fewer outliers in the neurons of the final layers are associated with enhanced zero-shot ICL performance. We visualize an example of such correlation in Figure 3. This observation implies that the kurtosis score of neuron activation distribution in FFN of the last decoder layer of LLMs could potentially serve as an effective signal for selecting optimal label names in zero-shot classification.
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast
Llama 3 (8b) Lowest Kurtosis 0.8431 0.7684 0.5516 0.6698 0.5870 0.6445 0.4097 0.3662 0.3190
Original Label 0.8187 0.5485 0.4387 0.6504 0.3416 0.5565 0.3487 0.3130 0.2395
Llama 3.1 (8b) Lowest Kurtosis 0.8789 0.8856 0.6212 0.7274 0.6403 0.6581 0.3642 0.3637 0.2458
Original Label 0.8064 0.6783 0.4426 0.6983 0.4337 0.6492 0.3528 0.3784 0.2126
Gemma 2 (9b) Lowest Kurtosis 0.9110 0.9247 0.6233 0.7595 0.5989 0.6784 0.4896 0.4858 0.3439
Original Label 0.8966 0.8728 0.6454 0.7707 0.5704 0.6874 0.4778 0.4858 0.3383
Flan-T5-xl (3b) Lowest Kurtosis 0.9334 0.9380 0.6881 0.5997 0.5813 0.4951 0.4759 0.4688 0.3857
Original Label 0.9305 0.8971 0.7267 0.6341 0.5580 0.5697 0.3837 0.4628 0.3473
Flan-T5-xxl (12b) Lowest Kurtosis 0.9428 0.9644 0.6905 0.6158 0.5938 0.5257 0.3852 0.6151 0.4218
Original Label 0.9407 0.9630 0.7814 0.7598 0.5614 0.5697 0.2016 0.6065 0.3278
Table 5: Comparison of zero-shot ICL performance on test sets between prompting with LOADS-selected label names versus original dataset label names. We underline the highest model performance (statistically significant with paired chi-squared test).
French Portuguese
Llama 3 (8b) Lowest Kurtosis 0.5020 0.4528
Original Label 0.4544 0.4284
Llama 3.1 (8b) Lowest Kurtosis 0.4728 0.4278
Original Label 0.3731 0.3679
Flan-T5-xl (3b) Lowest Kurtosis 0.5137 0.3912
Original Label 0.4189 0.3317
@@ 6.1 Method
    Choose the label set that leads to the lowest averaged kurtosis score.
6.2 Results
Evaluation
For evaluation, we compare the model performance of using label names selected by LOADS with the performance of using original label names from the dataset (i.e., the label names in Table 1), as using the original label name is the conventional and common practice when prompting the LLMs for classification tasks.
We randomly sample 100 data points from the validation set for label selection and test the selected label sets on the official test sets. Further, to examine whether our approach could generalize to other LLMs and LLMs with different model sizes, we also experiment with instruction-tuned Gemma-2-it (9b)  [^6] https://huggingface.co/google/gemma-2-9b-it and a larger version of Flan-T5, Flan-T5-xxl with 13 billion parameters.
Table 5 demonstrates that, on most of the datasets, employing LOADS to select label sets for zero-shot ICL prompts yields superior performance compared to directly using the original dataset label names. The improvement is consistent across diverse model architectures and sizes. [^7] We also test the prompt sensitivity in Appendix I.
@@ 6.2 Results
Previous research (Zhang et al., 2023a) has demonstrated that for non-English datasets, using English task instructions (including label names) while keeping the input in the original language often yields superior performance than non-English instructions. Therefore, we investigate whether the optimal English label sets selected by LOADS on English dataset can also enhance performance on non-English datasets in this scenario.
@@
Data Efficiency of LOADS
To show the merit of data efficiency of LOADS, we randomly sample 50, 100, 300, 500 or 1000 data points from the validation set and compare the rankings of the label sets based on LOADS. Due to the resource restriction, we conduct experiments on snopes and emergent datasets with Flan-T5-xl and Llama 3.
@@
Alternative Methods
We explore other approaches to improve LOADS, including prompt perplexity (Gonen et al., 2023) (see Appendix J) and model internal representation of label names (Wang et al., 2023) (see Appendix K), but they did not yield any significant correlation with zero-shot ICL model performance.
7 Related Work
ICL performance
Most work on few-shot ICL focuses on cases where LLMs are directly prompted with $N$ demonstration examples. The studies highlight the substantial impact of example ordering (Lu et al., 2022), formatting (Yoo et al., 2022; Wei et al., 2024; Mao et al., 2024; Liu et al., 2024a), and examples selection (Zhang et al., 2022; Liu et al., 2022a; Peng et al., 2024). Some work focuses on improving few-shot ICL performance via the optimal selection or arrangement of examples (Liu et al., 2022a; Rubin et al., 2022; Lu et al., 2022; Zhang et al., 2022; Liu et al., 2024b; Xu et al., 2024), re-weighting examples (Yang et al., 2023), automatic reformat or generation of demonstration representations (Kim et al., 2022; Liu et al., 2023), and introduction of intermediate reasoning steps (Wei et al., 2022; Zhang et al., 2023b). However, the impact of lexical choices for label names in classification tasks has received little attention, with the only closely related work suggesting that models are likely to confuse class categories, given their similar key vectors in the attention modules (Wang et al., 2023). For zero-shot ICL, Mu et al. (2024) demonstrate the effect of using synonyms for class options, but they neither consider the order of the label options nor propose an effective strategy to choose the label names. Gonen et al. (2023) empirically show that perplexity could be an effective indicator for prompt selection, but they do not account for class options. Yet different model behaviors between few-shot ICL and zero-shot ICL have been frequently observed, suggesting that findings from few-shot ICL do not necessarily hold in the zero-shot context (Lin and Lee, 2024).
Prompt-Tuning and Verbalizer
Prompt-tuning aims to automatically find or generate an optimal discrete prompt (e.g., through gradient-based search (Shin et al., 2020; Shi et al., 2023) or fine-tuning (Gao et al., 2021; Le Scao and Rush, 2021; Deng et al., 2022)) or by training continuous prompt (Lester et al., 2021; Liu et al., 2022b). Verbalizer is commonly followed after prompt-tuning or pattern exploit training (Gao et al., 2021) for the languages models pre-trained with masked language modelling. It aims to map output predictions to labels for classification tasks, including manually created verbalizer (Schick and Schütze, 2021), search-based verbalizer that identifies label words automatically from the dataset (Gao et al., 2021; Shin et al., 2020), and soft verbalizers that uses continuous embeddings obtained through fine-tuning (Hambardzumyan et al., 2021; Cui et al., 2022).
@@ 8 Conclusion
We study the impact of label options in the prompt for classification in zero-shot ICL, including lexical choice, labels’ order, and label elaborations. We observe a significant effect of the lexicons used to represent label names in the prompt, also linking to the impact of the labels’ order. Through neuron activation analysis, we find that optimal label sets produce fewer outlier neurons in LLMs’ feed-forward networks. We then propose LOADS, a novel method for selecting optimal label sets using activation distribution kurtosis. Prompting with label set selected by LOADS consistently outperforms the use of original dataset labels across different models. Our approach is post-hoc, data-efficient and requires no gradient propagation or model fine-tuning. It also demonstrates cross-lingual transferability when using English instructions for non-English datasets. By showing that carefully selecting label sets based on neuron activation patterns can significantly enhance model performance without requiring additional training or labeled data, this paper has important implications for leveraging LLMs in zero-shot classification scenarios.
Limitation
Our experiments focused primarily on stance classification tasks, and future studies should explore whether these findings generalize to a broader range of classification tasks and domains. Also, although we examined multiple models from the Flan-T5 and Llama families, our study did not include other popular language models (such as Phi [^8] https://huggingface.co/microsoft/phi-2 or Mistral  [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2) due to computational resource limitation. Expanding the range of models would provide a more comprehensive understanding of label option’s impact across different architectures.
@@
Our method, while more efficient than gradient-based approaches, still requires running inference on a subset of data to compute activation statistics. This may be challenging for resource-constrained environments or very large models. Also, we ensure the inclusion of samples for each class. The effectiveness of our method might vary when the validation set is highly imbalanced or even lack of data for the minority class. Also, the effectiveness of different distribution metrics is out of scope but we acknowledge that it may have significant improvement for our method.
@@
Ethics Statement
Our work has received approval from the Ethics Committee of our university and complies with X’s data policy for research.
Acknowledgement
This work is funded by EMIF managed by the Calouste Gulbenkian Foundation [^10] The sole responsibility for any content supported by the European Media and Information Fund lies with the author(s) and it may not necessarily reflect the positions of the EMIF and the Fund Partners, the Calouste Gulbenkian Foundation and the European University Institute. under the "Supporting Research into Media, Disinformation and Information Literacy Across Europe" call (ExU – project number: 291191).111111exuproject.sites.sheffield.ac.uk Yue Li is supported by a Sheffield–China Scholarships Council PhD Scholarship. The Portuguese and French translations of RumourEval were funded by the European Association for Machine Translation (EAMT). [^12] https://eamt.org We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing.
@@ Appendix A Datasets
We summarise the dataset characteristics in Table 7. For the datasets without official validation sets, we use the train/validation splits provided by Schiller et al. (2021).
Dataset Name Target Source Number of Label Sets
scd Topic Debates 31
perspectrum Claim Debates 31
snopes Claim News 31
ibmcs Topic Debates + Wikipedia 31
vast Topic Debates + Artificial 62
emergent Headline News 62
semeval Topic Social Media 93
rumoureval Tweet Social Media 248
arc Claim Debates 62
Table 7: Dataset characteristics (Hardalov et al., 2021) and the number of label sets we experiment with for each dataset.

### v2

@@ 4 Neuron Analysis for Label Selection
Although our findings indicate the importance of label word selection for text classification in zero-shot ICL, current studies lack consideration of this factor. Therefore, we conduct empirical analysis to gain insights into the underlying mechanism of lexical choice for single-word label names.
We preliminarily explored related approaches discussed in Section 2, including prompt perplexity Gonen et al. (2023) and model internal representation of label words Wang et al. (2023), but they did not yield any significant correlation with zero-shot ICL model performance (see the Appendix I and J for details). Meanwhile, various studies have indicated the correlation between neuron activation pattern in FFN and model performance Kuzmin et al. (2023); Tang et al. (2024); Stolfo et al. (2024); Wu et al. (2024). Inspired by the finding that the presence of outliers in neural networks is predictive of quantization and pruning performance for the layers of LLMs Kuzmin et al. (2023), we establish a new hypothesis: the model performance influenced by label names is correlated with the number of outliers in the neurons within FFN in the decoder of the LLMs. We empirically validate our hypothesis on the nine stance classification datasets, as well as two topic classification datasets (AG News Zhang et al. (2015) and TweetTopic Antypas et al. (2022)) to show the generalizability to other NLP tasks.
@@
where $\tilde{h}^{i}$ is the output hidden states from multi-head self-attention module. The activation function (act_fn) for Flan-T5 and Llama 3/3.1 is Gaussian Error Linear Unit (GELU) Hendrycks and Gimpel (2016) and Sigmoid Linear Unit (SiLU) Hendrycks and Gimpel (2016); Elfwing et al. (2018), respectively.
A neuron is defined as the linear transformation of each column in $W^{i}_{1}$ followed by the activation function. Here, we study the last layer $I$’s output of $\text{act\_fn}(\tilde{h}^{I}W^{I}_{1})$ (denoted as $N_{I}$) for the predicted first token of the label name in the model generation. Following Kuzmin et al., Kuzmin et al. (2023), we measure the number of outliers over the neuron output distribution ($N_{I}$) through kurtosis, given by:
@@
where $\mu$ is the mean of $N_{I}$. For each dataset, we average the kurtosis scores over the validation set for each candidate label set. We then calculate the Spearman correlation between model performance and the averaged kurtosis score.
@@ Results
Table 3 shows that, for most datasets, there is statistically significant negative correlation between model performances and kurtosis scores across models and activation functions, indicating that fewer outliers in the neurons of the final layers are associated with enhanced zero-shot ICL performance. This observation implies that the kurtosis score of neuron activation distribution in FFN of the last decoder layer of LLMs could potentially serve as an effective signal for selecting optimal label names in zero-shot classification.
Model Method Stance Classification Topic Classification
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast TweetTopic AG News
Llama 3 (8b) LOADS 0.8431 0.7684 0.5516 0.6698 0.5870 0.6445 0.4097 0.3662 0.3190 0.7752 0.7660
Original Label 0.8187 0.5485 0.4387 0.6504 0.3416 0.5565 0.3487 0.3130 0.2395 0.7742 0.7594
Original + Verbalizer 0.8185 0.5485 0.4306 0.6578 0.3400 0.5576 0.3476 0.3131 0.2395 0.7742 0.7594
Self-generated 0.6912 0.5802 0.3314 0.6607 0.3692 0.6032 0.2745 0.2837 0.3070 0.5851 0.6235
Llama 3.1 (8b) LOADS 0.8789 0.8856 0.6212 0.7274 0.6403 0.6581 0.3642 0.3637 0.2458 0.7988 0.7306
Original Label 0.8064 0.6783 0.4426 0.6983 0.4337 0.6492 0.3528 0.3784 0.2126 0.7909 0.7273
Original + Verbalizer 0.8064 0.6783 0.4426 0.6983 0.4262 0.6448 0.3534 0.3838 0.2128 0.7909 0.7273
Self-generated 0.6244 0.4918 0.5146 0.6798 0.2984 0.6421 0.3409 0.3373 0.2578 0.5956 0.6217
Gemma 2 (9b) LOADS 0.9110 0.9247 0.6233 0.7595 0.5989 0.6784 0.4896 0.4858 0.3439 0.8032 0.8451
Original Label 0.8966 0.8728 0.6454 0.7707 0.5704 0.6874 0.4778 0.4858 0.3383 0.8241 0.8316
Original + Verbalizer 0.8966 0.8728 0.6424 0.7707 0.5523 0.6873 0.4689 0.4748 0.3383 0.8246 0.8322
Self-generated 0.9017 0.9113 0.6656 0.7480 0.5833 0.6621 0.3549 0.4657 0.3178 0.6303 0.8215
Flan-T5-xl (3b) LOADS 0.9334 0.9380 0.6881 0.5997 0.5813 0.4951 0.4759 0.4688 0.3857 0.8530 0.8556
Original Label 0.9305 0.8971 0.7267 0.6341 0.5580 0.5697 0.3837 0.4628 0.3473 0.8071 0.9209
Original + Verbalizer 0.9305 0.8953 0.7026 0.6507 0.5586 0.5712 0.3852 0.4613 0.3482 0.8091 0.9209
Self-generated 0.7914 0.7384 0.3677 0.6974 0.3883 0.5954 0.2986 0.3343 0.3042 0.8501 0.7368
Flan-T5-xxl (13b) LOADS 0.9428 0.9644 0.6905 0.6158 0.5938 0.5257 0.3852 0.6151 0.4218 0.7727 0.7742
Original Label 0.9407 0.9630 0.7814 0.7598 0.5614 0.5697 0.2016 0.6065 0.3278 0.6643 0.9177
Original + Verbalizer 0.9407 0.9621 0.7716 0.7598 0.5631 0.5705 0.2011 0.6062 0.3278 0.6643 0.9177
Self-generated 0.8622 0.8384 0.4226 0.7315 0.4309 0.6182 0.3110 0.6115 0.2881 0.7242 0.9177
Table 4: Comparison of zero-shot ICL performance on test sets between prompting with LOADS-selected label names versus the other three baseline approaches. We underline the highest model performance (statistically significant with paired chi-squared test).
@@ 5.1 Method
    Choose the label set with the lowest averaged kurtosis score.
The above LOADS-selected label set can then be used in the standard zero-shot ICL on test sets.
5.2 Evaluation
Setups
We randomly sample 100 data points from the validation set for label selection and test the selected label sets on the official test sets.
  • Baselines: We compare the model performance of using label words selected by LOADS to the following three approaches: (1) Original label words: we use the original label words from the dataset (i.e., the labels in Table 1), as it is the conventional and widely adopted practice; (2) Original label words with a verbalizer: after prompting the LLMs with the original label words, we employ our pool of candidate label words (See Section 3.1) as a verbalizer and incorporate their probabilities into the final probability of the same class; (3) Self-generated label words: we prompt the LLMs without providing any class options and select the candidate label words with the average highest probability at the first generated label token.
  • LLMs: In addition to Llama3 (8b), Llama 3.1 (8b) and Flan-T5-xl (3b), we also examine whether LOADS could generalize to other model families and model sizes by including instruction-tuned Gemma-2-it (9b) and Flan-T5-xxl (13b).
Results
Table 4 presents the zero-shot ICL model performances on stance classification and topic classification datasets with different label word selection strategies. The results demonstrate that employing LOADS to select label sets for zero-shot ICL prompts yields superior performance compared to other baseline approaches on most of the datasets. The improvement is consistent across NLP tasks and datasets, model architectures and sizes, as well as prompt templates [^7] The analysis of prompt sensitivity can be found in the Appendix H..
Also, we observe limited benefits from adopting the verbalizer in post-processing, since LLMs tend to give the predicted label word tokens high probability in most of cases. Prompting with the self-generated label words rarely results in the best performance, while with the risk of leading to extremely low performance on certain datasets (e.g., perspectrum and snopes datasets with Llama 3).
Furthermore, potential data leakage could have significant impact on LOADS, indicated by the high performance achieved by the original label words on the AG News dataset with Flan-T5 models which was instruction-tuned with this dataset. It also aligns with the results in Table 3 where no statistically negative correlation was observed on the AG news dataset with Flan-T5-xl.
5.3 Analysis
@@
Previous research Zhang et al. (2023a) has demonstrated that for non-English datasets, prompting with English task instructions (including label words) while keeping the input in the original language often yields superior performance than non-English instructions. Therefore, we investigate whether the optimal English label sets selected by LOADS on English dataset can also enhance performance for non-English datasets in this scenario.
@@
Model Method French Portuguese
Llama 3 (8b) LOADS 0.5020 0.4528
Original Label 0.4544 0.4284
Llama 3.1 (8b) LOADS 0.4728 0.4278
Original Label 0.3731 0.3679
Flan-T5-xl (3b) LOADS 0.5137 0.3912
Original Label 0.4189 0.3317
Data Efficiency
To show the merit of data efficiency of LOADS, we randomly sample 50, 100, 300, 500 or 1000 data points from the validation set and compare the rankings of the label sets based on LOADS. Due to the resource restriction, we conduct experiments on snopes (binary classification) and emergent (three-way classification) datasets with Flan-T5-xl and Llama 3.
@@
Label Transferability
We explore whether LOADS-selected label sets can be generalised across datasets or even models for NLP tasks such as stance classification where different label lexicons are used to represent the same classes across datasets. Specifically, for each dataset $D_{i}$ on each LLM $M_{j}$, we select the optimal label set $L_{D_{i}M_{j}}$ through LOADS. To analyse whether the LOADS-selected label set on one dataset could be adapted to another related dataset with the same LLM $M$, we calculate the overlap of optimal labels ($L_{D_{i}\_M}$) between each dataset. Similarly, we examine the overlap of optimal labels ($L_{D\_M_{i}}$) between each model to explore whether the LOADS-selected optimal labels for dataset $D$ could be adapted across LLMs. We only focus on the positive and negative stances to enable comparison across binary, three-way and four-way stance classification datasets.
Our results indicate that LOADS-selected label sets is transferable across datasets on the same LLM, highlighting the potential of leveraging LOADS to identify optimal label words with established related datasets, avoiding the need to collect samples for the target new dataset. For example, the positive-negative stance label pairs identified for Llama 3 is endorse and deny across all the stance classification datasets. However, we find that the label words selected for a specific dataset on one LLM often differ from those identified for another LLM, suggesting LOADS’ dependency on the underlying model architectures and parameters.
In summary, the LOADS-selected label sets tend to be model-dependent rather than dataset-dependent. This observation aligns with the mechanism of LOADS, as the neurons and their distributions are inherently tied to the specific model. We hypothesize that this may suggest a correlation between the LOADS-selected label words and the LLMs’ internal representation or understanding of the target NLP task or concept (e.g., what is stance), highlighting potential directions for future studies.
@@ 6 Conclusion
We study the impact of label options in the prompt for classification in zero-shot ICL, including lexical choice, label order, and label elaborations. We observe a significant effect of the lexicons used to represent label words in the prompt, also linking to the models’ sensitivity to the label order. Through neuron activation analysis, we find that optimal label sets produce fewer outlier neurons in LLMs’ feed-forward networks. We then propose LOADS, a novel method for selecting optimal label sets using activation distribution kurtosis. Prompting with LOADS-selected label sets consistently outperforms the use of original dataset labels across different models. Our approach is post-hoc, data-efficient and requires no gradient propagation or model fine-tuning. It also demonstrates cross-lingual transferability when using English instructions for non-English datasets. By showing that carefully selecting label sets based on neuron activation patterns can significantly enhance model performance without requiring additional training or labeled data, this paper has important implications for leveraging LLMs in zero-shot classification.
Limitations
Our experiments focused primarily on stance classification tasks. We chose this task because the label ambiguity is an identified challenge (i.e., label names could be replaced by a sufficient number of synonyms without altering their meanings and scopes in the original study), and it has sufficient datasets for empirical study. Although we have tested the generalisabity of our findings on topic classification, with more datasets released and new tasks proposed in future studies, studies could be conducted to explore whether our findings generalize to a broader range of classification tasks and domains. Also, although we examined multiple models from the Flan-T5 and Llama families, our study did not include other popular language models (such as Phi [^8] https://huggingface.co/microsoft/phi-2 or Mistral  [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2) due to computational resource limitation. Expanding the range of models would provide a more comprehensive understanding of label option’s impact across different architectures.
@@
Our method, while more efficient than gradient-based approaches, still requires running inference on a subset of data to compute activation statistics. This may be challenging for resource-constrained environments or very large models. The efficiency of our method may also be challenged when the classification task contains a very large number of class categories. Furthermore, we ensure the inclusion of samples for each class. The effectiveness of our method might vary when the validation set is highly imbalanced or even lack of data for the minority class. The effectiveness of different distribution metrics is out of scope but we acknowledge that it may have significant improvement for our method.
@@
Acknowledgments
This work is funded by EMIF managed by the Calouste Gulbenkian Foundation [^10] The sole responsibility for any content supported by the European Media and Information Fund lies with the author(s) and it may not necessarily reflect the positions of the EMIF and the Fund Partners, the Calouste Gulbenkian Foundation and the European University Institute. under the "Supporting Research into Media, Disinformation and Information Literacy Across Europe" call (ExU – project number: 291191).111111exuproject.sites.sheffield.ac.uk and the UK’s innovation agency (InnovateUK) grant number 10039039 (approved under the Horizon Europe Programme as VIGILANT [^12] https://www.vigilantproject.eu, EU grant agreement number 101073921).
@@ Appendix A Datasets
We summarise the datasets we used in our study in Table 6. For the stance classification datasets without official validation sets, we use the train/validation splits provided by Schiller et al., Schiller et al. (2021). We utilize the official test set for AG news [^13] https://huggingface.co/datasets/sh0416/ag_news, and the official “train/test random split in the COLING 2022 paper" for TweetTopic dataset [^14] https://huggingface.co/datasets/cardiffnlp/tweet_topic_single. TweetTopic dataset has six class categories, potentially resulting in more than 4,000 different label sets if we consider only five synonymy words for each category (i.e., more than 12,000 experiments on three LLMs). Due to our limited computational resource, we experiment with three topics: pop culture, daily life, and science & technology.
Dataset Name Source # of Label Sets
scd Debates 31
perspectrum Debates 31
snopes News 31
ibmcs Debates + Wikipedia 31
vast Debates + Artificial 62
emergent News 62
semeval Social Media 93
rumoureval Social Media 248
arc Debates 62
AG News News 50
TweetTopic Social Media 64
Table 6: Datasets and the number of label sets we experiment with for each dataset.